OpenStudy (anonymous):

What is the value of n in the equation 8n + 6 = 10? Please enter your answer as a fraction.

OpenStudy (anonymous):

8n+6=10 Minus 6 from both sides 8n=4 divide 8 from both sides n=4/8 Simplify n=1/2

OpenStudy (a1234):

8n = 4 n = 4/8, or 1/2